Calculate the mole fraction of a solution containing 2.6 grams of sodium sulfate in 20 grams of water


1
Expert's answer
2022-03-09T09:06:03-0500

"n(Na_2SO_4)=2.6\ g\times\frac{1\ mol}{142.04\ g}=0.0183\ mol"


"n(H_2O)=20\ g\times\frac{1\ mol}{18.015\ g}=1.11\ mol"


"\chi(Na_2SO_4)=\frac{0.0183\ mol}{0.0183\ mol+1.11\ mol}=0.0162"


"\chi(H_2O)=\frac{1.11\ mol}{0.0183\ mol+1.11\ mol}=0.984"


Answer: Na2SO4 - 0.0162; H2O - 0.984